OpenStudy (darkigloo):

If the graph of y=f(x) contains the point (0,2), dy/dx=-x/y(e^(x^2)) and f(x)>0 for all x, the f(x)= A) 3+e^(-x^2) B)sqrt(3)+e^(-x) C)1+e^(-x) D)sqrt(3+e^(-x^2)) E)sqrt(3+e^(x^2))

zepdrix (zepdrix):

\[\Large\bf\sf y'\quad=\quad \frac{-x}{y e^{x^2}}, \qquad\qquad\qquad y(0)=2\]

zepdrix (zepdrix):

Ok so it appears our equation is separable.

zepdrix (zepdrix):

What part are you stuck on? The whole thing giving you trouble?

OpenStudy (darkigloo):

yes. i just don't know what to do.

zepdrix (zepdrix):

\[\Large\bf\sf \frac{dy}{dx}\quad=\quad \frac{-x}{y e^{x^2}}\]Multiply dx to the other side,\[\Large\bf\sf dy\quad=\quad \frac{-x}{y e^{x^2}}dx\]Multiply y over as well,\[\Large\bf\sf y\;dy\quad=\quad \frac{-x}{e^{x^2}}dx\]Rewrite your exponential using rules of exponents so it's easier to work with,\[\Large\bf\sf y\;dy\quad=\quad -x e^{-x^2}dx\]

zepdrix (zepdrix):

And from there, we can integrate each side.

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits y\;dy\quad=\quad \int\limits -x e^{-x^2}dx\]Do you understand how to integrate each side? The right side will require a u-substitution.

OpenStudy (darkigloo):

y^2/2=(-x^2/2)(e^-x^2)+c ?

zepdrix (zepdrix):

Oh oh hmm where is the x^2 coming from in front?

OpenStudy (darkigloo):

not sure what you mean..

zepdrix (zepdrix):

\[\Large\bf\sf \int\limits\limits -x e^{-x^2}dx\] \[\Large\bf\sf \mathcal{\text{Let}},\qquad u=-x^2\]\[\Large\bf\sf\mathcal{\text{Then,}}\qquad \frac{1}{2}du=-x\;dx\] \[\Large\bf\sf \int\limits\limits\limits\limits -x e^{-x^2}dx\quad=\quad \int\limits \frac{1}{2}e^{u}\;du\]Something like that, right? :o

zepdrix (zepdrix):

You should only end up with an exponential on the right side, not an x^2 in front. I don't understand where that came from.

zepdrix (zepdrix):

Oh you applied the power rule to the x and ignored the exponential? :O oh boy careful there!

OpenStudy (darkigloo):

ohh...ok. :(

OpenStudy (darkigloo):

sorry, can you explain how to integrate the whole thing.

zepdrix (zepdrix):

I thought I just ... did. +_+ You've already got the left side figured out. Do you not understand the substitution I made? Should I go through it slower?

OpenStudy (darkigloo):

i don't really get it. where did the x go in front of e^-x^2 and does int(ydy) become y^2/2?

zepdrix (zepdrix):

Lemme rearrange the right side, maybe it'll be easier to figure out what's going on this way. \[\Large\bf\sf \int\limits e^{\color{#DD4747}{-x^2}}\left(\color{royalblue}{-x\;dx}\right)\]Same problem, I've just grouped things a little differently so you can see how our substitution is being applied.

zepdrix (zepdrix):

\[\Large\bf\sf \color{#DD4747}{u=-x^2}\]

zepdrix (zepdrix):

Remember how to deal with a substitution? We take the derivative of u, what do you get? :O

OpenStudy (darkigloo):

derivative of -x^2 is -2x

zepdrix (zepdrix):

Looks good. \[\Large\bf\sf du=\color{royalblue}{-}2\color{royalblue}{x\;dx}\] This notation make sense? ^

zepdrix (zepdrix):

We want to solve for the blue part, -x dx. So we'll divide both sides by 2.

zepdrix (zepdrix):

\[\Large\bf\sf \color{royalblue}{\frac{1}{2}du=-x\;dx}\]

OpenStudy (darkigloo):

right. i get that

zepdrix (zepdrix):

So when we apply our substitution, \[\Large\bf\sf \int\limits\limits e^{\color{#DD4747}{-x^2}}\left(\color{royalblue}{-x\;dx}\right)\qquad\to\qquad \int\limits\limits e^{\color{#DD4747}{u}}\left(\color{royalblue}{\frac{1}{2}du}\right)\]

zepdrix (zepdrix):

It simplifies things down quite a bit,\[\Large\bf\sf \frac{1}{2}\int\limits e^u \;du\]

zepdrix (zepdrix):

Integrating gives:\[\Large\bf\sf \frac{1}{2}e^{u}+c\]

zepdrix (zepdrix):

You're still not understanding where the x went? :o

OpenStudy (darkigloo):

ohh now it see.

OpenStudy (darkigloo):

what do i do after that?

zepdrix (zepdrix):

Undo your substitution, giving the equation:\[\Large\bf\sf \frac{1}{2}y^2\quad=\quad \frac{1}{2}e^{-x^2}+c\]

zepdrix (zepdrix):

Multiply both sides by 2,\[\Large\bf\sf y^2\quad=\quad e^{-x^2}+C\]Any confusion with that step?

OpenStudy (darkigloo):

no, i understand.

zepdrix (zepdrix):

You can solve for y by taking the square root, buuuuuuuuuut... before we do that, we should use our initial data to solve for C. It will be easier to find C from this point.

OpenStudy (darkigloo):

c=3?

zepdrix (zepdrix):

Ok good!

OpenStudy (darkigloo):

yay. now what?

zepdrix (zepdrix):

All of your options a-e are in terms of `y=`. So we need to solve for y. Then you should be able to match it up to one of your options.

OpenStudy (darkigloo):

do i have to plug c in somewhere?

zepdrix (zepdrix):

\[\Large\bf\sf y^2\quad=\quad e^{-x^2}+3\]

zepdrix (zepdrix):

Yes, plug C in. And then set your y and x back to variables like that. ^

OpenStudy (darkigloo):

ohh so the answer is d?

zepdrix (zepdrix):

Yay good job!

OpenStudy (darkigloo):

Thank you so much! :D